OpenStudy (anonymous):

solve the system of equations by addition 3x -4y =5 x - 8y =20

OpenStudy (anonymous):

In general we want to cancel one of the variables when we are adding equations. Multiply the top here by -2. -6x+8y=-10 x-8y=20 now add -5x=10 x=-2 plug back in -2-8y=20 y=-11/4